# Bootstraping Clustering of Gaussians for View-consistent 3D Scene Understanding

### v1

@@ Conclusion
This paper presents FreeGS, a novel unsupervised method for 3D scene understanding. Based on the proposed IDSF and 2D-3D bootstrapping optimization strategy, the proposed method uniformly addresses a variety of tasks, including novel view segmentation, 3D detection and interactive object selection. However, some limitations remain in this label-free training paradigm. First, the introduced semantic fields bring extra memory and computational overhead to 3DGS, slightly impacting the model’s efficiency. Furthermore, there is still significant room for improvement in 3D detection performance. We believe that 3D tasks may represent the future direction for 3DGS development.

### v2

@@ Conclusion
In this paper, we present FreeGS, a novel unsupervised method for 3D scene understanding. Based on the proposed IDentity-coupled Semantic Field (IDSF) and 2D-3D collaborative bootstrapping optimization strategy, the proposed method eliminates the need for time-consuming preprocessing and uniformly addresses a variety of tasks, including novel view 2D segmentation, open vocabulary 3D detection and interactive object selection (as shown in Fig. 6).
However, some limitations remain in this label-free training paradigm. First, the introduced semantic fields bring extra memory and computational overhead to 3DGS, slightly impacting the model’s efficiency. Second, the Gaussian misalignment phenomenon and the large-scale low-texture background jointly result in suboptimal background segmentation. Furthermore, there is still significant room for improvement in 3D detection performance and we believe that 3D tasks may represent the future direction for 3DGS development.
Acknowledgments
This work was supported by National Natural Science Foundation of China under Grant 62206039, 62406053, 62293542, 62476048, and the Fundamental Research Funds for the Central Universities (DUT24RC(3)025).